# Helvipad: A Real-World Dataset for Omnidirectional Stereo Depth Estimation

### v1

@@ 5.4 Cross-Scene Generalization
To safely deploy learning systems in real-world settings, it is essential to assess their ability to generalize to unseen or underrepresented conditions, like outdoor environments at night. The bar chart in Fig. 5 shows the performance of our proposed omnidirectional adaptation, 360-IGEV-Stereo, alongside other evaluated methods, using depth MARE as the metric. We trained model variants on different subsets (Indoor, Outdoor, Indoor+Outdoor) and tested them under various conditions (Indoor, Outdoor, Night Outdoor). While models trained on specific environments perform well within their respective domains, indoor-trained models show a clear limitation in generalizing to outdoor scenes, particularly at night. Interestingly, this performance drop is less pronounced in omnidirectional models, such as 360-IGEV-Stereo and 360SD-Net, than in conventional models, like IGEV-Stereo and PSMNet. This results suggests that omnidirectional approaches benefit from improved cross-scene generalization capacities.
@@
Figure 6: Visualisation of disparity maps estimates using IGEV-Stereo and 360-IGEV methods against ground-truth sparse labels. The top row depicts images from the bottom camera; the second row provides the ground-truth sparse disparity maps; the third and fourth rows show the predicted disparity maps by IGEV-Stereo and 360-IGEV-Stereo respectively.
@@ 5.5 Qualitative Results
We provide illustrations to assess the qualitative results of IGEV-Stereo and 360-IGEV-Stereo methods. Visual comparisons between their respective predicted disparity maps and the ground truth are presented in Fig. 6. Overall, 360-IGEV-Stereo displays closer alignment to the ground truth. For example, in the right image, IGEV-Stereo fails to detect the two small people on the left, whereas 360-IGEV-Stereo successfully captures them, indicating a better capacity to capture small objects. In the left image, the use of circular padding by 360-IGEV-Stereo results in smoother transitions at the image boundaries. The incorporation of the polar angle also contributes to straighter lines and sharper visualizations, improving the representation of elements affected by distortions, such as pillars in the left image.
@@ 6 Conclusion
In this paper, we build Helvipad, a real-world stereo dataset for omnidirectional depth estimation, and demonstrate that incorporating polar angle as input and circular padding helps to adapt stereo matching deep learning models to equirectangular images. We also introduce a method for increasing depth label density from LiDAR point clouds through depth completion tailored to equirectangular projections. Despite achieving promising results, accurate depth estimation in omnidirectional contexts remains challenging. We hope this work encourages further research into omnidirectional stereo matching, especially for real-time applications. Moreover, this dataset is an ideal testbed for assessing the robustness of depth estimation methods to diverse lighting conditions and depth ranges.
@@ Acknowledgments
This work was supported by the EPFL Center for Imaging through a Collaborative Imaging Grant. We thank the VITA lab members for their valuable feedback, which helped to enhance the quality of this manuscript. We also express our gratitude to Dr. Simone Schaub-Meyer and Oliver Hahn for their insightful advice during the project’s final stages.
@@ A.1 Sequences
The Helvipad dataset includes 29 video sequences captured between December 2023 and February 2024., recorded at a frame rate of 10 Hz.
@@ B.1 Data Capture Setup
The hardware setup consists of a custom-designed support system that aligns all devices horizontally and stacks them vertically. The system integrates two Ricoh Theta V cameras, which capture images in 4K/UHD equirectangular format at 30 fps with an initial resolution of 3840 × 1920 pixels. It also includes an Ouster OS1-64 LiDAR sensor, operating at 10 fp with 64 beams and a vertical field of view of 42.4°. The LiDAR is mounted at the bottom, with the first camera at the top (“top camera”) and the second camera in the middle (“bottom camera”). The vertical distances between the devices are precisely configured, with 19.1 cm separating the two camera lenses and 45 cm between the LiDAR and the bottom camera, as shown in Fig. 7.
@@ C.1 Evaluation Method
In Sec. 3.3, the ratio of interpolated points (RIP) is defined as the ratio of interpolated points after filtering. As the uncertainty estimates of all query points within one sequence do not fit into memory, thresholds for the uncertainty-based filtering are calculated for each point cloud and then averaged per sequence. This leads to an actual ratio of interpolated points (ARIP) which differs slightly from the desired RIP.
@@ C.2 Choice of Hyperparameters
Figure 14: Selected metrics with a variable number of aggregated point clouds (Fig. 14(a) and Fig. 14(b)) or $m=4$ previous / next point clouds (Fig. 14(c) and Fig. 14(d)) on all, indoor, outdoor and night outdoor train sequences. The ratio of interpolated points is set to $\text{RIP}=80\%$ and $k=4$ is chosen as the numbers of neighbors (Fig. 14(a) and Fig. 14(b)) or a variable number of neighbors $k$ (Fig. 14(c) and Fig. 14(d)).
@@ C.3 Temporal Aggregation Comparison
Table 6: Evaluation metrics for different temporal aggregation methods with $m=1$, $\text{RIP}=0.8$ and $k=4$ on all test sequences.
@@ C.4 Results
Table 7: Evaluation metrics for the final hyperparameters of the depth completion on all, indoor, outdoor, and night outdoor train sequences.
@@
Tab. 8 exhibits that the ratio of labeled points (RLP), as defined in Sec. C.1, is increased approximately by a factor of 5 following the application of depth completion across all sequence types. The maximum number of labelable pixels $\text{n}_{\text{lab, max}}$ corresponds to the denominator in Eq. 8.
@@ D.2 Implementation Details
We convert the disparity, given in degree, to pixels to be able to warp the appropriately for constructing the cost volume:
@@
During training 360-IGEV-Stereo had some instabilities. To mitigate this issue, we clamp the disparity $d_{\text{deg}}$ in each step to $d_{\text{deg}}\in[d_{\text{deg, min}},d_{\text{deg, max}}]$. According to the statistics of the dataset, the minimum and maximum disparity are set to $d_{\text{deg, min}}=0.062^{\circ}$ and $d_{\text{deg, max}}=14.75^{\circ}$.
To enable a better understanding of the context, we use the the full image size of 512x1920 for training. Except of the photometric data augmentation of the original code implementation, we do not apply any data augmentations.
All weights that have not been modified are initialized with the original IGEV-Stereo weights created with pretraining on Sceneflow by the authors. The model is trained with a batch size of 4 for 20 epochs which corresponds to around 130k steps. Furthermore, the maximum disparity for constructing the cost volumes is set to 96 which is the smallest number that is divisible by 32 and larger than the maximum disparity in pixels.
@@ 3. Multi-scale outputs
Method Initialization Disparity (°) Depth (m)
PSMNet [6] random 0.29 0.49 0.23 2.49 5.62 0.19
Scene Flow 0.33 0.54 0.29 2.78 6.17 0.19
360SD-Net [28] random 0.21 0.42 0.18 2.14 5.12 0.15
Stereo-MP3D 0.23 0.44 0.20 2.31 5.41 0.16
IGEV-Stereo [32] random 0.23 0.44 0.18 2.10 5.30 0.17
Scene Flow 0.22 0.41 0.17 1.85 4.44 0.15
360-IGEV-Stereo random 0.20 0.40 0.16 1.91 4.60 0.14
Scene Flow 0.18 0.39 0.15 1.77 4.36 0.14
Method Initialization Disparity (°) Depth (m)
360SD-Net [28] reported [28] 0.14 0.69 0.06 0.22
random 0.15 0.99 0.07 0.09 0.29 0.05
Helvipad 0.13 0.93 0.06 0.07 0.25 0.04
Table 13: Fine-tuning results on Stereo-MP3D [28] dataset.
@@ E.1 Effect of Pretrained Weights
E.2 Cross-Dataset Generalization
We study the cross-dataset generalization of 360SD-Net by first training the model on the Helvipad dataset and then fine-tuning it on the Stereo-MP3D [28] dataset, which shares a similar top-bottom camera configuration. Results are available in Tab. 13. Compared to a baseline trained from random initialization on Stereo-MP3D, fine-tuning on Helvipad significantly improved all depth and disparity metrics. For instance, the depth MAE decreased from 0.09 m to 0.07 m. Note that although we used the authors’ provided code [^5] https://github.com/albert100121/360SD-Net and the reported hyperparameters, our reproduced results differ from those reported in the original paper. We present our outcomes for a fair comparison.

### v2

@@ 5.4 Cross-Scene Generalization
To safely deploy learning systems in real-world settings, it is essential to assess their ability to generalize to unseen or underrepresented conditions, like outdoor environments at night. The bar chart in Fig. 5 shows the performance of our proposed omnidirectional adaptation, 360-IGEV-Stereo, alongside other evaluated methods, using depth MARE as the metric. We trained model variants on different subsets (Indoor, Outdoor, Indoor+Outdoor) and tested them under various conditions (Indoor, Outdoor, Night Outdoor). While models trained on specific environments perform well within their respective domains, indoor-trained models show a clear limitation in generalizing to outdoor scenes, particularly at night. Interestingly, this performance drop is less pronounced in omnidirectional models, such as 360-IGEV-Stereo and 360SD-Net, than in conventional models, like IGEV-Stereo and PSMNet. This result suggests that omnidirectional approaches benefit from improved cross-scene generalization capacities.
@@ 5.5 Qualitative Results
We present visualizations of the depth completion pipeline and qualitative results of IGEV-Stereo and 360-IGEV-Stereo on test images in Fig. 6. Depth completion produces dense, coherent disparity maps aligned with the ground truth and does not incorporate incorrect labels at depth boundaries. For stereo matching, 360-IGEV-Stereo displays closer alignment to the ground truth. For example, in the right image, IGEV-Stereo misses two small people on the left, whereas 360-IGEV-Stereo successfully captures them, indicating a better capacity to detect small objects. In the left image, the use of circular padding by 360-IGEV-Stereo results in smoother transitions at the image boundaries. The incorporation of the polar angle also contributes to straighter lines and sharper visualizations, improving the representation of elements affected by distortions, such as pillars in the left image.
@@ 6 Conclusion
We build Helvipad, a real-world stereo dataset for omnidirectional depth estimation, and show that incorporating polar angle as input and circular padding help to adapt deep stereo matching models to equirectangular images. We also introduce a depth completion pipeline tailored to equirectangular projections which increases label density from LiDAR point clouds. We hope this work encourages further research into omnidirectional stereo matching, especially for real-time applications. This dataset is an ideal testbed for assessing the robustness of depth estimation methods to diverse lighting conditions and depth ranges.
@@ Acknowledgments
This work was supported by the EPFL Center for Imaging through a Collaborative Imaging Grant. We thank the VITA lab members for their valuable feedback, which helped to enhance the quality of this manuscript. We also express our gratitude to Dr. Simone Schaub-Meyer and Oliver Hahn for their insightful advice during the project’s final stages. Additionally, we sincerely thank the anonymous reviewers for their constructive feedback and thoughtful suggestions, which significantly contributed to improving the clarity and robustness of this work.
@@ A.1 Sequences
The Helvipad dataset includes 26 video sequences captured between December 2023 and February 2024., recorded at a frame rate of 10 Hz.
@@ B.1 Data Capture Setup
The hardware setup consists of a custom-designed support system that aligns all devices horizontally and stacks them vertically. The system integrates two Ricoh Theta V cameras, which capture images in 4K/UHD equirectangular format at 30 fps with an initial resolution of 3840 × 1920 pixels. It also includes an Ouster OS1-64 LiDAR sensor, operating at 10 fp with 64 beams and a vertical field of view of 42.4°. The LiDAR is mounted at the bottom, with the first camera at the top (“top camera") and the second camera in the middle (“bottom camera"). The vertical distances between the devices are precisely configured, with 19.1 cm separating the two camera lenses and 45 cm between the LiDAR and the bottom camera, as shown in Fig. 7.
@@ C.1 Evaluation Method
In Sec. 3.3, the RIP is defined as the ratio of interpolated points after filtering. As the uncertainty estimates of all query points within one sequence do not fit into memory, thresholds for the uncertainty-based filtering are calculated for each point cloud and then averaged per sequence. This leads to an ARIP which differs slightly from the desired RIP.
@@ C.2 Choice of Hyperparameters
Figure 14: Selected metrics with a variable number of aggregated point clouds (Fig. 14(a) and Fig. 14(b)) or $m=4$ previous / next point clouds (Fig. 14(c) and Fig. 14(d)) on all, indoor, outdoor and night outdoor train sequences. The ratio of interpolated points is set to $\text{RIP}=80\%$ and $k=4$ is chosen as the numbers of neighbors (Fig. 14(a) and Fig. 14(b)) or a variable number of neighbors $k$ (Fig. 14(c) and Fig. 14(d)). We report all depth metrics in m
@@ C.3 Temporal Aggregation Comparison
Table 6: Evaluation metrics for different temporal aggregation methods with $m=1$, $\text{RIP}=0.8$ and $k=4$ on all test sequences. We report all depth metrics in m. The best results are highlighted in bold.
@@ C.4 Results
Table 7: Evaluation metrics for the final hyperparameters of the depth completion on all, indoor, outdoor, and night outdoor train sequences. We report all depth metrics in m.
@@
Tab. 8 exhibits that the RLP, as defined in Sec. C.1, is increased approximately by a factor of 5 following the application of depth completion across all sequence types. The maximum number of labelable pixels $\text{n}_{\text{lab, max}}$ corresponds to the denominator in Eq. 7.
@@ D.1 Evaluation Metrics
  • Left-Right Consistency Error (LRCE): This metrics evaluates the consistency at the left and right boundaries of 360° images by measuring the discrepancy between predicted depth values across the image edges. In the original work introducing LRCE [30], the metrics also accounts for left-right discrepancies in ground-truth data to address extreme cases where object edges align exactly with the image boundaries. Due to the sparsity of the LiDAR depth maps in the test set, there are very few valid points simultaneously at both image edges for computing LRCE metric (3 pixel pairs per image in average). As an alternative, we use the depth-completed tests maps (136 pixels in average) and compute LRCE with this augmented ground-truth. Given $\mathcal{B}_{i}$ the subset of valid pixel pairs in image $i$ where ground-truth labels exist for both the leftmost and rightmost columns ($\mathcal{B}_{i}\subset\mathcal{A}_{i}$), LRCE is defined as the sum of absolute differences between left and right edges for both predicted and ground-truth disparity values:
    $\text{LRCE}=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}\frac{1}{|\mathcal{B}i|}\sum_{j\in\mathcal{B}i}\left|e^{\text{gt}}_{i,j}-e^{\text{pred}}_{i,j}\right|,$ (16)
    where $e_{i,j}=\left|d_{\text{left},i,j}-d_{\text{right},i,j}\right|$ is the left-right discrepancy term in image $i$ for pixel pair $j$, computed for predicted disparity error ($e^{\text{pred}}_{i,j}$) and ground-truth disparity error ($e^{\text{gt}}_{i,j}$).
@@ D.2 Implementation Details
We convert the disparity, given in degree, to pixels to be able to warp the top image appropriately for constructing the cost volume:
@@
During training 360-IGEV-Stereo had some instabilities. To mitigate this issue, we clamp the disparity $d_{\text{deg}}$ in each step to $d_{\text{deg}}\in[d_{\text{deg, min}},d_{\text{deg, max}}]$. According to the statistics of the dataset, the minimum and maximum disparity are set to $d_{\text{deg, min}}=0.048^{\circ}$ and $d_{\text{deg, max}}=23^{\circ}$.
To enable a better understanding of the context, we use the the full image size of 512 x 1920 for training. Except of a common photometric data augmentation, we do not apply any data augmentations.
All weights that have not been modified are initialized with the original IGEV-Stereo weights created with pretraining on Sceneflow by the authors. The model is trained with a batch size of 4 for 20 epochs which corresponds to around 130k steps. Furthermore, the maximum disparity for constructing the cost volumes is set to 128 which is the smallest number that is divisible by 32 and larger than the maximum disparity in pixels.
@@ E.1 Effect of Pretrained Weights
E.2 Left-Right Consistency
In addition to the main results available in Tab. 2, we provide a more detailed analysis of Left-Right Consistency Error (LRCE) across different scene types in Tab. 12.
Model All Indoor Outdoor Night Outdoor
PSMNet 1.80 0.93 1.31 1.16
360SD-Net 0.90 0.52 1.02 1.01
IGEV-Stereo 1.20 0.79 1.21 1.55
360-IGEV-Stereo $\bm{0.38}$ $\bm{0.17}$ $\bm{0.38}$ $\bm{0.46}$
Table 12: Depth-LRCE with augmented ground-truth across different scene types. Results are reported in meters.
The results indicate that left-right consistency is more challenging to maintain in outdoor scenes, with the highest errors observed in night outdoor conditions due to low-light environments and reduced texture details. In contrast, indoor scenes achieve the lowest LRCE, likely due to more structured environments with well-defined depth boundaries and fewer extreme lighting variations.
E.3 Cross-Dataset Generalization
We study the cross-dataset generalization of 360SD-Net by first training the model on the Helvipad dataset and then fine-tuning it on Stereo-MP3D and Stereo-SF3D datasets [34] , which share a similar top-bottom camera configuration. Results are available in Tab. 14. Compared to a baseline trained from random initialization on Stereo-MP3D, fine-tuning on Helvipad significantly improved all depth and disparity metrics. On Stereo-MP3D, fine-tuning on Helvipad reduces depth MAE from 0.087m (random initialization) to 0.072m, along with consistent improvements in all disparity metrics. A similar trend is observed on Stereo-SF3D, where Helvipad pretraining improves depth MAE from 0.029m to 0.027m and disparity MAE from 0.105° to 0.099°, outperforming models pretrained on Stereo-MP3D. Note that although we used the authors’ provided code [^5] https://github.com/albert100121/360SD-Net and the reported hyperparameters, our reproduced results differ from those reported in the original paper. We present our outcomes for a fair comparison.
@@ E.3 Cross-Dataset Generalization
Method Initialization Disparity (°) Depth (m)
PSMNet [6] random 0.29 0.50 0.25 2.51 5.67 0.18
Scene Flow 0.33 0.54 0.29 2.78 6.17 0.19
360SD-Net [34] random 0.22 0.42 0.19 2.12 5.08 0.18
Stereo-MP3D 0.23 0.44 0.20 2.31 5.41 0.16
IGEV-Stereo [38] random 0.23 0.44 0.18 2.10 5.30 0.17
Scene Flow 0.23 0.42 0.17 1.86 4.47 0.15
360-IGEV-Stereo random 0.20 0.40 0.16 1.91 4.60 0.14
Scene Flow 0.19 0.40 0.15 1.72 4.30 0.13
Dataset Initialization Disparity (°) Depth (m)
Stereo-MP3D reported [34] 0.145 0.693 0.059 0.218
random 0.148 0.994 0.074 0.087 0.294 0.050
Helvipad $\bm{0.129}$ $\bm{0.930}$ $\bm{0.063}$ $\bm{0.072}$ $\bm{0.252}$ $\bm{0.039}$
Stereo-SF3D reported [34] 0.103 0.369 0.003 0.091
random 0.105 $\bm{0.468}$ 0.020 0.029 0.071 0.016
Stereo-MP3D 0.121 0.505 0.023 0.035 0.079 0.019
Helvipad $\bm{0.099}$ 0.469 $\bm{0.018}$ $\bm{0.027}$ $\bm{0.069}$ $\bm{0.015}$
Table 14: Cross-dataset generalization results by fine-tuning 360SD-Net [34].
Real-world representation. While collected on a university campus, the dataset captures many common urban environments, such as parking lots, roads, underpasses, pedestrian squares, footpaths, corridors and crowded halls. To further demonstrate transferability, we present below the qualitative result of 360-IGEV-Stereo on a real-world image without labels from the 360SD-Net paper:
Figure 18: Qualitative result of 360-IGEV-Stereo on a real-world image. The top image shows the bottom image of the input, while the bottom image displays the predicted disparity map.
Trained solely on Helvipad, the model demonstrates zero-shot capabilities in this environment.